OpenStudy (anonymous):

equation of the tangent line of y=(1/2)arccosx points are posted in!

OpenStudy (anonymous):

\[(\frac{ -\sqrt{2} }{ 2 }, \frac{ 3\Pi }{ 8 })\]

OpenStudy (anonymous):

Okay, so first you find the derivative of \(y\).

OpenStudy (anonymous):

\[\frac{dy}{dx}=-\frac{1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

Remember the tangent line will have the form\[y=mx+b\]We use the derivative to find \(m\).

OpenStudy (anonymous):

Then we plug the coordinate into the x and y and we can solve for b.